# Pair Production in time-dependent Electric field at Finite times

### v4

@@ III Pair Production in Sauter-pulse electric field
$A\left(t\right)=-E_{0}\tau\textrm{tanh}\left(\frac{t}{\tau}\right).$ (45)
@@
Now, in the presence of external electric field Eq. (44) equation of motion for mode function Eq. (14) reads
$\Bigl{(}\partial_{t}^{2}+\textrm{i}E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)+\omega^{2}(\bm{p},t)\Bigr{)}\psi_{\bm{p}}(t)=0.$ (46)
where we have skipped the irrelevant index $\lambda$ of the mode function $\psi_{\bm{p}}(t).$ This equation is solved analytically by converting it into a hypergeometric differential equation. By changing the time variable to
$y=\frac{1}{2}\left(1+\textrm{tanh}\left(\frac{t}{\tau}\right)\right).$ (47)
@@
$\partial_{t}=(\partial_{t}y)\partial_{y}=\frac{2}{\tau}y(1-y)\partial_{y}$ (48)
$\partial_{t}^{2}=\frac{4}{\tau^{2}}y(1-y)\partial_{y}y(1-y)$ (49)
Now, Eq.(46) can be rewritten as
@@
with, $k=\frac{-\textrm{i}\tau\omega_{0}}{2},$ $l=\frac{\textrm{i}\tau\omega_{1}}{2},$ $\omega_{0}=\sqrt{\epsilon_{\perp}^{2}+P_{0}^{2}},$ $\omega_{1}=\sqrt{\epsilon_{\perp}^{2}+P_{1}^{2}},$ $P_{0}=\left(p_{\parallel}-eE_{0}\tau\right),$ $P_{1}=\left(p_{\parallel}+eE_{0}\tau\right),$
in Eq. (50), we have $\eta_{\bm{p}}(y)$ satisfying the following hypergeometric differential equation Abramowitz and Stegun (1964)
$\left(y\left(1-y\right)\partial_{y}^{2}+\left(c-\left(a+b+1\right)y\right)\partial_{y}-ab\right)\eta_{\bm{p}}(y)=0.$ (52)
@@
$a$ $=-\textrm{i}E_{0}\tau^{2}-\frac{\textrm{i}\tau\omega_{0}}{2}+\frac{\textrm{i}\tau\omega_{1}}{2}=i\zeta_{1}$
@@
$\eta_{\bm{p}}^{(-)}(y)=y^{1-c}(1-y)^{c-a-b}{}_{2}\mathcal{F}_{1}\left(1-a,1-b,2-c;y\right).$ (55)
@@
To get the mode functions $\psi_{\bm{p}}^{(\pm)}(y)$ we have to resubstitute $\eta_{\bm{p}}^{(\pm)}(y)$ in our ansatz Eq.(51)
@@
where $N^{(\pm)}(\bm{p})$ as normalization constants (see section A.1 for detailed calculations).
@@ III.1 Particle distribution function
To obtain the single-particle distribution function using the function $\beta(\bm{p},t)$ in terms of the new time variable $y$. This transformation yields
@@
$f(\bm{p},y)$ $=|N^{+}(\bm{p})|^{2}\Biggl{(}\frac{\omega(\bm{p},y)+P(p_{\parallel},y)}{\omega(\bm{p},y)}\Biggr{)}\bigg{|}\frac{2}{\tau}y(1-y)\frac{ab}{c}f_{1}+\textrm{i}(\omega(\bm{p},y)-(1-y)\omega_{0}-y\omega_{1})f_{2}\bigg{|}^{2},$ (59)
where $f_{1}={}_{2}\mathcal{F}_{1}\left(1+a,1+b,1+c;y\right)$, $f_{2}={}_{2}\mathcal{F}_{1}\left(a,b,c;y\right).$ This expression for the time-dependent single-particle distribution function gives information about the pair production at the various dynamical stagesBlaschke et al. (2011a); Banerjee and Singh (2022). In this context, we focus on analyzing the evolution of the distribution function from the initial QEPP stage to the final REPP stage, passing through the transient stage. Within this formalism, the distribution function encodes the spectral information of produced particles, and its behavior is examined in detail in Sec. IV. Additionally, in Sec. III.3, we derive a simple approximate expression for the distribution function, valid for $t>>\tau$.
Furthermore, in order to analyze the characteristics of the vacuum polarization effect during the formation of pairs, we derive vacuum polarization and depolarization functions using the analytical expression of Bogoliubov coefficients in the proceeding section III.2.
@@ III.2 Correlation function
As can easily be seen, this function $\mathcal{C}(\bm{p},t)$, consisting of creation operators for a particle and an anti-particle with the opposite momentum describes the process of production of $e^{-}e^{+}$ pair. In several research articles Schmidt et al. (1998); Banerjee and Singh (2019); Blaschke et al. (2017); Fedotov et al. (2011); Banerjee and Singh (2022), the particle-antiparticle correlation function is redefined by incorporating the slowly varying component of the time-dependent creation and annihilation operators in adiabatic number basis.
^$\hat{\mathcal{B}}_{{\bm{p}}r}(t)=\hat{B}_{{\bm{p}}r}(t)\textrm{e}^{-\textrm{i}\Theta_{\bm{p}}(t)},$ (62)
$\hat{\mathcal{D}}_{-{\bm{p}}r}(t)=\hat{D}_{-{\bm{p}}r}(t)\textrm{e}^{-\textrm{i}\Theta_{\bm{p}}(t)}.$ (63)
@@
$=\textrm{e}^{2\textrm{i}\Theta_{\bm{p}}(t)}\langle{0_{in}}|\hat{D}_{-{\bm{p}}r}^{\dagger}(t)\hat{B}_{{\bm{p}}r}^{\dagger}(t)|{0_{in}}\rangle,$
$=2\alpha_{\bm{p}}^{*}(t)\beta_{\bm{p}}(t)\textrm{e}^{2\textrm{i}\Theta_{\bm{p}}(t)}.$ (64)
Now, using the above relation for the pair correlation function in terms of the transformed time variable $y$, we proceed similarly to the derivation of the single-particle distribution function $f(\bm{p},t)$.
@@
$\mathcal{C}(\bm{p},t)$ $=|N^{+}(\bm{p})|^{2}\frac{\epsilon^{2}_{\perp}(p_{\perp})}{\omega(\bm{p},y)}\Biggl{[}\frac{4}{\tau^{2}}y^{2}(1-y)^{2}|\frac{ab}{c}|^{2}|f_{1}|^{2}-(\omega^{2}(\bm{p},y)-(y\omega_{1}+(1-y)\omega_{0})^{2})|f_{2}|^{2}$
$-2y(1-y)(\omega_{1}-\omega_{0}-2E_{0}\tau)\left((y\omega_{1}+(1-y)\omega_{0})\Re{(\frac{b}{c}f_{1}f^{*}_{2})}+\textrm{i}\omega(\bm{p},y)\Im{(\frac{b}{c}f_{1}f^{*}_{2})}\right)\Biggr{]}$ (65)
As we know, the vacuum polarization effects play a crucial role in the process of pair production. This effect is defined through the functions $u(\bm{p},t)=\Re{(\mathcal{C}(\bm{p},t))}$ and $v(\bm{p},t)=\Im{(\mathcal{C}(\bm{p},t))}$Banerjee and Singh (2019).
$u(\bm{p},t)$ $=|N^{+}(\bm{p})|^{2}\frac{\epsilon^{2}_{\perp}(p_{\perp})}{\omega(\bm{p},y)}\Biggl{(}\frac{4}{\tau^{2}}y^{2}(1-y)^{2}|\frac{ab}{c}|^{2}|f_{1}|^{2}-(\omega^{2}(\bm{p},y)-(y\omega_{1}+(1-y)\omega_{0})^{2})|f_{2}|^{2}$
$-2y(1-y)(\omega_{1}-\omega_{0}-2E_{0}\tau)(y\omega_{1}+(1-y)\omega_{0})\Re{(\frac{b}{c}f_{1}f^{*}_{2})}\Biggr{)}$ (66)
$v(\bm{p},t)$ $=|N^{+}(\bm{p})|^{2}2\textrm{i}y(1-y)(2E_{0}\tau+\omega_{0}-\omega_{1})\Im{(\frac{b}{c}f_{1}f^{*}_{2})}$ (67)
@@
The function $u(\bm{p},t)$ depicts vacuum polarization effects and pair production phenomena. The function $v(\bm{p},t)$ serves as counterterm to pair production, effectively representing pair annihilation in the vacuum excitation process. A comprehensive discussion of these insights is presented in detail in Sec. IV.3.
III.3 Approximate analytical expression for $f(\bm{p},t)$ at finite time
@@
To investigate the behavior of the distribution functions in the limit $t>\tau,$ we employ approximations based on the suitable expressions of the Gamma and the Gauss-hypergeometric functions. These approximations enable us to deduce a simplified analytical expression for particle distribution functions. First, we start with approximating the Gauss-hypergeometric function as $y\to 1$. It is crucial to ensure smooth convergence towards the limit of ${}_{2}\mathcal{F}_{1}\left(a,b,c;y\rightarrow 1\right)$. This task is complicated by the intricate nature of the parameters $a$, $b$, and $c$ in this specific context, which makes it essential to exercise caution when dealing with this limit. Therefore, it is beneficial to transform the argument by substituting $y$ with $(1-y)$. This transition can be achieved using the following mathematical identity.
${}_{2}\mathcal{F}_{1}\left(a,b,c;z\right)=\frac{\Gamma\left(c\right)\Gamma\left(c-a-b\right)}{\Gamma\left(c-a\right)\Gamma\left(c-b\right)}{}_{2}\mathcal{F}_{1}\left(a,b,a+b-c+1;1-z\right)\\ +\left(1-z\right)^{c-a-b}\frac{\Gamma\left(c\right)\Gamma\left(a+b-c\right)}{\Gamma\left(a\right)\Gamma\left(b\right)}{}_{2}\mathcal{F}_{1}\left(c-a,c-b,c-a-b+1;1-z\right).\\ \left\lvert\textrm{arg}(1-z)\right\rvert<\pi$ (68)
In general Gauss-hypergeometric function,
@@
Using the above relations Eq. (68) and  (69), we approximate the Gauss-hypergeometric functions $f_{1}$ and $f_{2}$ present in the relation of the particle distribution function (see Eq.(59)) as follows:
@@
Here, $y$ is a variable that changes with time according to Eq. (47). As time progresses, especially in the case with significant time intervals such as $t>\tau$, it becomes evident that $(1-y)$ approaches zero. Consequently, the dominant contribution to the particle distribution function as $y\rightarrow 1$ originates from the zeroth-order term, which is independent of $(1-y).$ As discussed in many literature for deriving the asymptotic expression for the particle distribution function, the Gauss-hypergeometric functions are truncated up to zeroth order onlyHebenstreit et al. (2010); Gavrilov and Gitman (1996).
Building on the earlier discussion (see Sec. III.1), our focus is primarily on understanding the dynamics of particle distribution function over finite time rather than just its asymptotic trends. We want to examine the distribution function, $f(\bm{p},y)$, in the vicinity of $y\rightarrow 1.$ In that case, it is necessary to incorporate other higher-order terms in the expression of the asymptotic particle distribution function. It means expanding the time-dependent particle distribution function defined in Eq. (59) as a power series in the variable $(1-y)$ as:
$f(\bm{p},y)\simeq\mathrm{C}_{0}(\bm{p},y)+(1-y)\mathrm{C}_{1}(\bm{p},y)+(1-y)^{2}\mathrm{C}_{2}(\bm{p},y)+...+(1-y)^{n}\mathrm{C}_{n}(\bm{p},y).$ (72)
To compute an approximate expression for the particle distribution function that depends on finite time, we can truncate the power series of the Gauss-hypergeometric functions $f_{1}$ and $f_{2}$ given by Eqs. (70) and (71) respectively, up to a specific order. The order of truncation will depend on the desired accuracy and the characteristics of the finite-time behavior under consideration.

### v5

@@ III Pair Production in Sauter-pulse electric field
$A\left(t\right)=-E_{0}\tau\textrm{tanh}\left(\frac{t}{\tau}\right),$ (46)
@@
Now, in the presence of an external Sauter-pulse field, the equation of motion for the mode function Eq. (14) reads
$\Bigl{(}\partial_{t}^{2}+\textrm{i}E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)+\omega^{2}(\bm{p},t)\Bigr{)}\psi_{\bm{p}}(t)=0,$ (47)
where we have skipped the irrelevant index $\lambda,$ of the mode function $\psi_{\bm{p}}(t).$ This equation is solved analytically by converting it into a hypergeometric differential equation. By changing the time variable to
$y$ $=\frac{1}{2}\left(1+\textrm{tanh}\left(\frac{t}{\tau}\right)\right),\quad\text{with}\quad y\in[0,1],$ (48)
so that
$P(p_{\parallel},y)$ $=p_{\parallel}+eE_{0}\tau(2y-1),$ (49)
$\omega^{2}(p_{\parallel},y)$ $=\epsilon^{2}(p_{\perp})+P^{2}(p_{\parallel},y).$ (50)
@@
$\partial_{t}=(\partial_{t}y)\partial_{y}=\frac{2}{\tau}y(1-y)\partial_{y},$ (51)
$\partial_{t}^{2}=\frac{4}{\tau^{2}}y(1-y)\partial_{y}y(1-y).$ (52)
Equation (47) can be rewritten as follows
@@
where
$k=\frac{-\textrm{i}\tau\omega_{0}}{2},\quad l=\frac{\textrm{i}\tau\omega_{1}}{2},$
and we define
$P_{0}=P(p_{\parallel},y=0)=p_{\parallel}-eE_{0}\tau,\quad P_{1}=P(p_{\parallel},y=1)=p_{\parallel}+eE_{0}\tau,$ (55)
$\omega_{0}=\omega(\bm{p},y=0)=\sqrt{\epsilon_{\perp}^{2}+P_{0}^{2}},\quad\omega_{1}=\omega(\bm{p},y=1)=\sqrt{\epsilon_{\perp}^{2}+P_{1}^{2}}.$ (56)
Inserting this ansatz into Eq. (53), we have $\eta_{\bm{p}}(y)$ satisfying the following hypergeometric differential equation Abramowitz and Stegun (1964)
$\left(y\left(1-y\right)\partial_{y}^{2}+\left(c-\left(a+b+1\right)y\right)\partial_{y}-ab\right)\eta_{\bm{p}}(y)=0,$ (57)
@@
$a$ $=-\textrm{i}E_{0}\tau^{2}-\frac{\textrm{i}\tau\omega_{0}}{2}+\frac{\textrm{i}\tau\omega_{1}}{2}=i\zeta_{1},$
@@
$\eta_{\bm{p}}^{(-)}(y)=y^{1-c}(1-y)^{c-a-b}{}_{2}\mathcal{F}_{1}\left(1-a,1-b,2-c;y\right),$ (60)
@@
To get the mode functions $\psi_{\bm{p}}^{(\pm)}(y)$ we have to resubstitute $\eta_{\bm{p}}^{(\pm)}(y)$ in the ansatz Eq.(54)
@@
with $N^{(\pm)}(\bm{p})$ being normalization factors.
We chose the normalization factors such that the mode functions $\psi^{(+)}_{\bm{p}}(y)$ and the functions $\phi_{p}^{(\pm)}(y)$ coincide for $y\rightarrow 0:$
$\psi^{(\pm)}_{\bm{p}}(y)=\phi^{(\pm)}_{\bm{p}}(y)\quad\text{for}\quad y\rightarrow 0.$ (63)
The asymptotic behavior of the mode function for $y\rightarrow 0$ is given by:
$\psi^{(+)}_{\bm{p}}(y)\stackrel{{\scriptstyle y\to 0}}{{=}}N^{(+)}(\bm{p})e^{k\ln{(y)}},$ (64)
$\psi^{(-)}_{\bm{p}}(y)\stackrel{{\scriptstyle y\to 0}}{{=}}N^{(-)}(\bm{p})e^{-k\ln{(y)}}.$ (65)
The asymptotic behavior of the adiabatic mode functions:
$\phi_{\bm{p}}^{(\pm)}(y)=\frac{e^{\mp\textrm{i}\Theta_{\bm{p}}(y_{0},y)}}{\sqrt{2\omega(\bm{p},y)\bigl{(}\omega(\bm{p},y)\mp P(p_{\parallel},y)\bigr{)}}},$ (66)
with
[…]
is determined in order to fix the value of normalization constants $N^{(\pm)}(\bm{p}).$
One needs to integrate the phase carefully, since the $1/y$ factor leads to a divergence at $y=0$. We follow the limiting integration procedure described in Ref. Hebenstreit et al. (2010). The dynamical phase $\Theta_{\bm{p}}(y_{0},y)$ diverges as $y\rightarrow 0$. To isolate this divergence, the phase can be split into two parts:
$\Theta_{\bm{p}}(y_{0},y)\stackrel{{\scriptstyle y\to 0}}{{=}}\frac{\tau}{2}\omega_{0}\ln{y}+\tilde{\Theta}_{\bm{p}}(y_{0},y=0).$ (68)
The first term captures the divergent behavior, while $\tilde{\Theta}_{\bm{p}}(y_{0},y)$ represents the regular, finite part. This regular term depends on a reference point $y_{0}\not=\{0,1\}$, which is introduced to regulate the lower limit of the integral in the phase expression.
So that the asymptotic behavior of the adiabatic mode functions $\phi_{\bm{p}}^{(\pm)}(y)$ reads
$\phi_{\bm{p}}^{(\pm)}(y)=\frac{e^{\mp\textrm{i}\tilde{\Theta}_{\bm{p}}(y_{0},0)}}{\sqrt{2\omega_{0}\bigl{(}\omega_{0}\mp P_{0}\bigr{)}}}e^{\mp\textrm{i}\frac{\tau}{2}\omega_{0}\ln{(y)}}.$ (69)
The normalization constants $N^{(\pm)}(\bm{p})$ are accordingly given by:
$N^{(\pm)}(\bm{p})=\frac{e^{\mp\textrm{i}\tilde{\Theta}_{\bm{p}}(y_{0},0)}}{\sqrt{2\omega_{0}\bigl{(}\omega_{0}\mp P_{0}\bigr{)}}}.$ (70)
@@ III.1 Particle distribution function
To obtain the single-particle distribution function using the Bogoliubov coefficient $\beta(\bm{p},t)$ in terms of the new time variable $y$. This transformation yields
@@
$f(\bm{p},y)$ $=|N^{(+)}(\bm{p})|^{2}\Biggl{(}1+\frac{P(p_{\parallel},y)}{\omega(\bm{p},y)}\Biggr{)}\bigg{|}\frac{2}{\tau}y(1-y)\frac{ab}{c}f_{1}+\textrm{i}\biggl{(}\omega(\bm{p},y)-(1-y)\omega_{0}-y\omega_{1}\biggr{)}f_{2}\bigg{|}^{2},$ (72)
where $|N^{(+)}(\bm{p})|^{2}=\frac{1}{{2\omega_{0}\bigl{(}\omega_{0}\mp P_{0}\bigr{)}}},f_{1}={}_{2}\mathcal{F}_{1}\left(1+a,1+b,1+c;y\right)$, and $f_{2}={}_{2}\mathcal{F}_{1}\left(a,b,c;y\right).$
This expression for the time-dependent single-particle distribution function provides insight into pair production at various dynamical stages Blaschke et al. (2011a); Banerjee and Singh (2022). In this context, we focus on analyzing the evolution of the distribution function from the initial QEPP stage to the final REPP stage, passing through the transient stage Banerjee and Singh (2022). Within this framework, the distribution function encodes the spectral information of the produced particles, and its behavior is discussed in Sec. IV. Additionally, in Sec. III.3, we derive a simple approximate expression for the distribution function, valid for $t\gg\tau$.
It is important to note that $y$ is a transformed time variable used to obtain the analytic solution of the differential equation. This, in turn, leads to an analytical relation for the distribution function, which depends on time $t$. However, we employ the $y$-variable as time to derive an approximate relation for the distribution function. Plots of the momentum distribution function, or momentum spectrum, will be shown for different time values ($t$-values), which will allow us to compare our results with those from the literature wherever possible.
@@ III.2 Correlation function
As can easily be seen, this function $\mathcal{C}(\bm{p},t)$, consisting of creation operators for a particle and an anti-particle with the opposite momentum, describes the process of production of $e^{-}e^{+}$ pair.
In several research articles Schmidt et al. (1998); Banerjee and Singh (2019); Blaschke et al. (2017); Fedotov et al. (2011); Banerjee and Singh (2022), the particle-antiparticle correlation function is redefined by incorporating the slowly varying component of the time-dependent creation and annihilation operators in adiabatic number basis:
^$\hat{\mathcal{B}}_{{\bm{p}}r}(t)=\hat{B}_{{\bm{p}}r}(t)\textrm{e}^{-\textrm{i}\Theta_{\bm{p}}(t_{0},t)},$ (75)
$\hat{\mathcal{D}}_{-{\bm{p}}r}(t)=\hat{D}_{-{\bm{p}}r}(t)\textrm{e}^{-\textrm{i}\Theta_{\bm{p}}(t_{0},t)}.$ (76)
@@
$=\textrm{e}^{2\textrm{i}\Theta_{\bm{p}}(t_{0},t)}\langle{0_{in}}|\hat{D}_{-{\bm{p}}r}^{\dagger}(t)\hat{B}_{{\bm{p}}r}^{\dagger}(t)|{0_{in}}\rangle,$
$=2\alpha_{\bm{p}}^{*}(t)\beta_{\bm{p}}(t)\textrm{e}^{2\textrm{i}\Theta_{\bm{p}}(t_{0},t)}.$ (77)
Using the above relation, the pair correlation function can be calculated using the Bogoliubov coefficients (Eq.(39),(40)) and the mode function $\psi_{\bm{p}}^{(+)}(y)$( Eq. (61):
@@
$\mathcal{C}(\bm{p},t)$ $=|N^{(+)}(\bm{p})|^{2}\frac{\epsilon^{2}_{\perp}(p_{\perp})}{\omega(\bm{p},y)}\Biggl{[}\frac{4}{\tau^{2}}y^{2}(1-y)^{2}\left|\frac{ab}{c}f_{1}\right|^{2}-\Bigl{(}\omega^{2}(\bm{p},y)-(y\omega_{1}+(1-y)\omega_{0})^{2}\Bigr{)}|f_{2}|^{2}$
$-\left(2y(1-y)(\omega_{1}-\omega_{0}-2E_{0}\tau)\right)\left((y\omega_{1}+(1-y)\omega_{0})\Re{(\frac{b}{c}f_{1}f^{*}_{2})}+\textrm{i}\omega(\bm{p},y)\Im{(\frac{b}{c}f_{1}f^{*}_{2})}\right)\Biggr{]}.$ (78)
Vacuum polarization effects play a fundamental role in the process of pair production, influencing the dynamics of particle creation in strong fields. These effects are captured by the correlation function $\mathcal{C}(\bm{p},t)$, which can be decomposed into its real and imaginary parts to describe different aspects of the process  Banerjee and Singh (2019); Aleksandrov:2020mez; Smolyansky:2019dqd. Specifically, the real part of $\mathcal{C}(\bm{p},t)$ is denoted as $u(\bm{p},t)=\Re\left(\mathcal{C}(\bm{p},t)\right)$, while the imaginary part is given by $v(\bm{p},t)=\Im\left(\mathcal{C}(\bm{p},t)\right)$. These components are crucial for understanding the underlying mechanisms of pair production and its dependence on time and momentum.
Therefore,
$u(\bm{p},t)$ $=|N^{(+)}(\bm{p})|^{2}\frac{\epsilon^{2}_{\perp}(p_{\perp})}{\omega(\bm{p},y)}\Biggl{(}\frac{4}{\tau^{2}}y^{2}(1-y)^{2}\left|\frac{ab}{c}f_{1}\right|^{2}-(\omega^{2}(\bm{p},y)-(y\omega_{1}+(1-y)\omega_{0})^{2})|f_{2}|^{2}$
$-2y(1-y)(\omega_{1}-\omega_{0}-2E_{0}\tau)(y\omega_{1}+(1-y)\omega_{0})\Re{(\frac{b}{c}f_{1}f^{*}_{2})}\Biggr{)},$ (79)
$v(\bm{p},t)$ $=2\textrm{i}y(1-y)\epsilon^{2}_{\perp}(p_{\perp})|N^{(+)}(\bm{p})|^{2}(2E_{0}\tau+\omega_{0}-\omega_{1})\Im{(\frac{b}{c}f_{1}f^{*}_{2})}.$ (80)
@@
The function $u(\bm{p},t)$ depicts vacuum polarization effects and pair production phenomena. The function $v(\bm{p},t)$ serves as a counterterm to pair production, effectively representing pair annihilation in the vacuum excitation process Aleksandrov:2020mez. A comprehensive discussion of these insights is presented in detail in Sec. IV.3.
III.3 Approximate analytical expression for the distribution function at finite time
To investigate the behavior of the distribution functions in the asymptotic limit $t\gg\tau$, we analyze the governing expressions for $f(\bm{p},y)$, which involve the Gamma and Gauss-hypergeometric functions. In this limit, we employ suitable asymptotic approximations for these special functions, as given in standard references Abramowitz and Stegun (1964). These approximations allow us to derive a simplified analytical expression for the particle distribution functions, capturing their leading behavior at late times.
@@ III.3 Approximate analytical expression for the distribution function at finite time
First, we start with approximating the Gauss-hypergeometric function as $y\to 1$. It is crucial to ensure smooth convergence towards the limit of ${}_{2}\mathcal{F}_{1}\left(a,b,c;y\rightarrow 1\right)$. This task is complicated by the intricate nature of the parameters $a$, $b$, and $c$ in this specific context, which makes it essential to exercise caution when dealing with this limit. Therefore, it is beneficial to transform the argument by substituting $y$ with $(1-y)$. This transition can be achieved using the following mathematical identity
${}_{2}\mathcal{F}_{1}\left(a,b,c;z\right)=\frac{\Gamma\left(c\right)\Gamma\left(c-a-b\right)}{\Gamma\left(c-a\right)\Gamma\left(c-b\right)}{}_{2}\mathcal{F}_{1}\left(a,b,a+b-c+1;1-z\right)\\ +\left(1-z\right)^{c-a-b}\frac{\Gamma\left(c\right)\Gamma\left(a+b-c\right)}{\Gamma\left(a\right)\Gamma\left(b\right)}{}_{2}\mathcal{F}_{1}\left(c-a,c-b,c-a-b+1;1-z\right).$ (81)
where $\left\lvert\textrm{arg}(1-z)\right\rvert<\pi$ for the identity to hold Abramowitz and Stegun (1964).
In general, the Gauss-hypergeometric function can be expressed as a power series:
@@
The series continues with additional terms involving higher powers of $z.$ Each term in the series involves the parameters $a,b,$ and $c$ as well as the variable $z$ raised to a specific power Abramowitz and Stegun (1964).
Using the above relations Eq. (81) and  (82), we approximate the Gauss-hypergeometric functions $f_{1}$ and $f_{2}$ present in the Eq.(72) as follows:
@@
$+(1-y)^{(c-a-b)}\frac{\Gamma(c)\Gamma(a+b-c)}{\Gamma(a)\Gamma(b)}\Biggl{(}1+\frac{(c-a)(c-b)}{(1+c-a-b)}(1-y)+...\Biggr{)}.$ (84)
Note that here $y$ is a transformed time variable that evolves according to Eq. (48).
As time progresses, especially in the case with significant time intervals such as $t>\tau$, it becomes evident that $(1-y)$ approaches zero. Consequently, the dominant contribution to the particle distribution function as $y\rightarrow 1$ originates from the zeroth-order term, which is independent of $(1-y).$ As discussed in many literature for deriving the asymptotic expression for the particle distribution function, the Gauss-hypergeometric functions are truncated up to zeroth order onlyHebenstreit et al. (2010); Gavrilov and Gitman (1996).
Building on the earlier discussion (see Sec. III.1), our focus is primarily on understanding the dynamics of the particle distribution function over finite time rather than just its asymptotic trends. We want to examine the distribution function, $f(\bm{p},y)$, in the vicinity of $y\rightarrow 1.$ In that case, it is necessary to incorporate other higher-order terms in the expression of the asymptotic particle distribution function. It means expanding the time-dependent distribution function defined in Eq. (72) as a power series in the variable $(1-y)$ as:
$f(\bm{p},y)\approx\mathrm{C}_{0}(\bm{p},y)+(1-y)\mathrm{C}_{1}(\bm{p},y)+(1-y)^{2}\mathrm{C}_{2}(\bm{p},y)+...+(1-y)^{n}\mathrm{C}_{n}(\bm{p},y).$ (85)
To compute an approximate expression for the distribution function that depends on finite time, we can truncate the power series of the Gauss-hypergeometric functions $f_{1}$ and $f_{2}$ given by Eqs. (83) and (84) respectively, up to a specific order. The order of truncation will depend on the desired accuracy and the characteristics of the finite-time behavior under consideration.
@@
Additionally, we can approximate
$1+\frac{P(p_{\parallel},y)}{\omega(\bm{p},y)}$ $\approx\Omega_{0}+(1-y)\Omega_{1}+(1-y)^{2}\Omega_{2},$ (90)
with
$\Omega_{0}$ $=1+\frac{P_{1}}{\omega_{1}},$ (91)
$\Omega_{1}$ $=\frac{-2E_{0}\tau}{\omega_{1}^{3}}\epsilon^{2}_{\perp}(p_{\perp}),$ (92)
We also approximate the cross term $\Gamma_{1}\overline{\Gamma_{2}}$, which appears in the expression for the distribution function. The detailed calculation of this term using the approximation for the Gamma function is provided in Appendix A.1. Here, we simply write final expression for the $\Gamma_{1}\overline{\Gamma_{2}}$ :
$\Gamma_{1}\overline{\Gamma_{2}}$ $\approx\left|\Gamma_{1}\overline{\Gamma_{2}}\right|e^{\textrm{i}\varrho},$ (94)
where the modulus and phase are given by
$\left|\Gamma_{1}\overline{\Gamma_{2}}\right|$ $=\frac{\omega_{0}}{2\omega_{1}}\cdot\frac{\exp\left[\frac{\pi\tau}{2}(\omega_{0}-\omega_{1}+2E_{0}\tau)\right]}{\sinh(\pi\tau\omega_{0})\sqrt{1+\omega_{1}^{2}\tau^{2}}}\sqrt{\frac{(\omega_{0}+2E_{0}\tau)^{2}-\omega_{1}^{2}}{\omega_{1}^{2}-(2E_{0}\tau-\omega_{0})^{2}}},$ (95)
and
$\varrho$ $=\frac{\tau}{2}(\omega_{0}+\omega_{1}-2E_{0}\tau)\ln\left[\tau(\omega_{0}+\omega_{1}-2E_{0}\tau)\right]+(-\omega_{0}+\omega_{1}-2E_{0}\tau)\ln\left[\tau(\omega_{0}-\omega_{1}+2E_{0}\tau)\right]$
$\quad+(-\omega_{0}+\omega_{1}+2E_{0}\tau)\ln\left[\tau(-\omega_{0}+\omega_{1}+2E_{0}\tau)\right]+(\omega_{0}+\omega_{1}+2E_{0}\tau)\ln\left[\tau(\omega_{0}+\omega_{1}+2E_{0}\tau)\right]$
$\quad+\pi-\tan^{-1}(\tau\omega_{1})-2\tau\omega_{1}\ln(2\omega_{1}\tau).$ (96)